### v2

# thanks: This study was supported by the Russian Science Foundation (Project No. 23-62-10026) [RSF]

Abstract
We present beam-dynamics simulations for an S-band RF photoinjector test bench under development at JINR, aimed at producing high-quality electron beams and enabling future generation of relativistic vortex electrons. Simulations of the 1.5-cell photogun under the presently available RF gradient of $\sim 45$ MV/m demonstrate stable bunch formation at low charge ($Q=0.63$ pC), where space-charge forces are weak and the transverse emittance is dominated by RF-induced correlations. Optimization of the injection phase and cathode-region solenoid yields a robust emittance-stabilization regime with a final normalized emittance of $\sim 3.2\ \pi\cdot\mathrm{mm\cdot mrad}$. To assess prospects for accelerating structured electron beams, we additionally model the quantum evolution of single-electron Laguerre–Gaussian wave packets. The results show that multi-MeV acceleration suppresses free-space spreading and preserves the wave-packet’s initial OAM structure, indicating that the test bench provides suitable conditions for forthcoming experimental studies of relativistic vortex electrons.
preprint: APS/123-QED
I Introduction
The generation of relativistic vortex electron beams – quantum states with a helical phase front and a quantized orbital angular momentum projection (OAM) – is attracting growing interest across accelerator, atomic, and nuclear physics, electron microscopy, quantum optics, and foundation of quantum mechanics [Knyazev18, Bliokh17, Karlovets21, Ivanov22]. To date, vortex electrons have been generated only in the sub-MeV domain, up to kinetic energies of $\sim 300$ keV of transmission electron microscopes [Uchida10, Verbeeck10, Blackburn14]. Extending such structured states into the multi-MeV regime relevant for photoinjectors and linacs remains an outstanding challenge, yet promises significant impact, including new tools for nuclear and hadronic structure studies and potential alternatives to spin-polarized beams [Ivanov22].
It is important to distinguish the genuine vortex beams of single particles whose wave functions are characterized by a helical phase structure $\exp(i\ell\varphi)$ and a quantized OAM projection $\ell\hbar$ from beams with modified transverse emittance ratios – such as flat and round beams demonstrated at Fermilab [Danilov96, Brinkmann02]. These latter beams can posses a so-called extrinsic angular momentum, which is a property of the beam as a whole, so a quantum state of each particle in the beam is not defined, whereas the former quantum beams of single electrons posses the so-called intrinsic angular momentum, the property of each electron akin to spin [Bliokh17, Floettmann20, Karlovets21, Ivanov22].
One feasible route towards high-energy vortex electrons is the transfer of OAM from twisted ultraviolet light to electrons during photoemission [Karlovets21, Pavlov24, Kazinski25]. Whereas the generation of pure vortex states with a definite angular momentum projection and its vanishing quantum uncertainty is challenging due the finite temperature, space-charge effects, and other factors [Floettmann20, Karlovets21, Kazinski25], the generation of a superposition of vortex states with an uncertainty less than the mean OAM value of the beam seems feasible when the OAM is very large, $\ell\gg\hbar$, and the current is low so the space-charge effects can be mitigated. Realizing this mechanism requires a photoinjector capable of producing stable, low-emittance beams with well-controlled transverse profiles and rapid acceleration, ensuring that the initial OAM-imprinted phase structure of the emitted electron is preserved during the early acceleration.. To this end, a dedicated S-band RF photoinjector test bench is being developed at the Laboratory of Nuclear Problems, Joint Institute for Nuclear Research (JINR), combining a 1.5-cell RF gun [Nikiforov18] and a high-power UV laser system [Gacheva14]. Along with relativistic vortex electrons, there are plans at the Institute of Modern Physics in China to put a twist on relativistic ions [An25].
RF photoinjectors are widely used as high-brightness electron sources at facilities such as LCLS [Alley99], the European XFEL [Ferrario01], and SuperKEKB [Cahill14], owing to their ability to quickly suppress space-charge effects via high-gradient acceleration [Giribono23, Maltseva24]. At JINR, there exists a LINAC facility, which is a linear accelerator of electrons to the energy of 200 MeV and plans to come to 400 MeV in 2027 [Nozdrin25], it operates with a thermionic DC gun and a multi-stage RF bunching system [Nozdrin21], producing $\sim 1$22 ps bunches. Replacing it with an RF photoinjector enables direct generation of $\sim 10$ ps bunches with the significantly improved transverse beam quality and precise control over the charge and spot size – critical prerequisites for producing structured quantum states.
In this work, we present beam-dynamics simulations of the bunch generation and acceleration in the JINR photoinjector using realistic field maps. We demonstrate the implementation of a Carlsten-type emittance-compensation scheme: although the emittance does not decrease by an order of magnitude, the solenoidal lens stabilizes the transverse emittance at a low value and suppresses RF-induced phase-space distortions, establishing a robust operating point for subsequent vortex-beam experiments.
We additionally investigate the quantum evolution of the single-electron Laguerre–Gaussian (LG) wave packets with different OAM values, up to $\ell=64\hbar$, accelerated in the same realistic DC and RF fields. The single-electron regime is justified by the planned experimental capability of operating at ultralow bunch charges approaching the quasi–single-electron limit. Collective effects such as Coulomb repulsion and decoherence are not included in this first study; thus the results represent an upper bound on coherence and OAM-structure preservation during acceleration.
II Test Bench Design
The RF photoinjector test bench under construction at JINR is based on an S-band 1.5-cell RF photogun and a high-power UV drive laser. The overall layout (Fig. 1) includes the RF gun, the laser beamline, two solenoids for transverse beam control, and a 1.5-m vacuum transport section operating at $\sim 10^{-9}\ \mathrm{Torr}$.
Figure 1: Schematic layout of the RF photoinjector test bench. Key components include the S-band RF photogun, the UV drive laser, the solenoid lenses, the vacuum laser beamline, and the laser-injection mirror.
The high accelerating field rapidly increases the electron energy, helping preserve beam quality at the earliest stage. Two solenoids – a cathode solenoid near the gun and a downstream focusing solenoid – shape the transverse phase space and control the beam size.
The 1.5-cell S-band RF gun operates in the $\pi$-mode at 2856 MHz [Samofalova21]. Its internal geometry and photocathode placement (Fig. 2) are optimized for rapid extraction and emittance-compensating beam formation. The UV laser enters through the coaxial channel, which simultaneously serves as the RF coupling element and the electron-beam exit port. The on-axis accelerating field (Fig. 3) was obtained from CST Microwave Studio eigenmode simulations [CST].
Figure 2: Internal geometry of the S-band radiofrequency photogun. The design consists of 1.5 accelerating cells operating in the $\pi$-mode at 2856 MHz. The central aperture serves both for laser injection and electron beam extraction.
Figure 3: Simulated axial distribution of the accelerating electric field along the beam propagation axis $z$, obtained using eigenmode analysis in CST Microwave Studio.
Although the gun is designed for up to 6MW6\ $\mathrm${MW} of input RF power, the present test-bench operation is limited to about 3MW3\ $\mathrm${MW} because the cavity is still undergoing RF conditioning. Thus the $\approx 45\ \mathrm{MV/m}$ accelerating gradient used in the simulations corresponds to the experimentally achievable field at this commissioning stage.
[…]
The emittance-compensation solenoid prototype (Fig. 4) was manufactured at JINR. It has an inner radius of 100 mm, outer radius of 140 mm, length of 100 mm, and contains 216 copper turns in six layers.
Figure 4: Mechanical layout and positioning of the emittance compensation solenoid along the beamline.
The axial field profile, measured with 5-mm resolution at 20 A, is shown in Fig. 5. These measured values are used directly in the beam-dynamics simulations.
Figure 5: Measured axial magnetic field distribution of the emittance compensation solenoid. The measurement was carried out in 5 mm steps along the $z$-axis using a magnetic field meter with a coil current of 20 A.
III Beam Dynamics Simulation
All beam dynamics simulations were carried out with ASTRA [ASTRA], which couples a 3D particle-in-cell space-charge solver (in cylindrical symmetry) to high-order Runge–Kutta tracking through RF-cavity and magnetic fields. The key inputs are the bunch charge (from photocathode quantum efficiency and laser energy), mean initial energy and its spread, initial normalized emittance and spatial/momentum distributions, and the RF phase at photoemission.
The initial beam parameters are derived from the properties of the UV drive laser and the copper photocathode. The fourth-harmonic laser operates at $\lambda=262\ \mathrm{nm}$, corresponding to a photon energy of 4.74eV4.74\ $\mathrm${eV}. With typical UV pulse energies of $3~\mu\mathrm{J}$ delivered to the vacuum chamber and taking into account the measured transport losses, the number of photons incident on the cathode is of order $4\times 10^{11}$ per pulse.
Using a copper photocathode with a quantum efficiency of $QE\sim 10^{-5}$ [Xiang15], this yields an emitted charge of approximately
Q=0.63pC,Q=0.63\ $\mathrm${pC},
which is used as the baseline value in the simulations.
The photoemission model follows the Dowell–Schmerge approach [Dowell09], assuming Fermi–Dirac statistics and Gaussian spatial distributions. For an effective copper work function of $\phi_{\rm eff}=4.3\ \mathrm{eV}$ [Karkare15], the resulting mean kinetic energy and energy spread of emitted electrons lie in the sub-eV range, consistent with standard metallic photocathodes. With the UV laser spot size set to $\sigma_{x,y}=2\ \mathrm{mm}$ (equal to the initial RMS beam size), the corresponding intrinsic normalized emittance is
$\varepsilon_{n,\,x,y}=1.07\ \pi\cdot\mathrm{mm\cdot mrad},$
which defines the initial phase-space conditions used for the beam-dynamics simulations.
A simulated RF-phase scan was performed to determine the optimal injection phase. Figure 6(a) shows the mean beam energy as a function of the initial RF phase $\varphi_{0}$. Maximum acceleration is achieved near $\varphi_{0}\approx 180^{\circ}$, with a phase tolerance of about $\pm 1^{\circ}$.
The phase scan of the normalized transverse emittance (Fig. 6(b)) shows a pronounced overall increase with injection phase in the explored interval. This behaviour is primarily determined by the intrinsic field structure of the 1.5-cell S-band RF gun operating in the $\pi$-mode: the axial electric field has a node close to the cathode plane, so the extraction field falls rapidly as the phase approaches the zero crossing. At the present low bunch charge (0.63pC0.63\ $\mathrm${pC}) the emittance is thus governed mainly by RF-induced transverse correlations imposed by the cavity geometry and mode structure, rather than by space-charge forces. This field configuration explains the observed monotonic phase dependence of the emittance. The operational working point of the test bench, chosen to maximize the energy gain under the available RF gradient, lies at $\varphi_{0}\approx 185^{\circ}$$190^{\circ}$. In this region the field on the cathode is not maximal, which inevitably increases the extraction-induced emittance, but it provides the highest attainable beam energy at the current stage of commissioning. The results of this scan define the initial parameters used in the subsequent beam-dynamics simulations (Table 1).
Figure 6: Simulated RF-phase scan results: (a) mean beam energy and (b) final normalized transverse emittance versus injection phase $\varphi_{0}$. Maximum acceleration occurs near $\varphi_{0}\approx 180^{\circ}$, while the operational working region $\varphi_{0}\approx 185^{\circ}$$190^{\circ}$ reflects the compromise between emittance and achievable energy.
Table 1: Initial parameters for beam dynamics simulations.
Parameter Value
Photon energy 4.74 eV
Work function (Cu) 4.3 eV
Mean kinetic energy 0.29 eV
RMS energy spread 0.10 eV
RMS beam size at cathode 2.0 mm
Normalized emittance 1.07  $\pi\cdot$ $m$ $\cdot$ mrad
Number of emitted electrons $3.95\times 10^{6}$
Bunch charge 0.63 pC
Optimal RF phase $185^{\circ}$
Total beamline length 1.568 m
III.1 Beam Dynamics Without Emittance Compensation
This process is identical for both simulation scenarios, with and without the emittance-compensating solenoid, as it depends solely on the amplitude of the RF field and the proper injection phase, and is independent of external magnetic fields.
Fig.7 shows the simulated mean kinetic energy of the bunch along the RF photogun beamline. A two‐peak structure arises from the axial field in the S‐band RF photogun, with rapid acceleration to $\overline{W_{\rm kin}}=1.9$ MeV at the gun exit, followed by a plateau. High‐gradient injection quickly suppresses space‐charge driven emittance growth once $\gamma\gg 1$, although early nonlinear distortions remain significant. Longitudinally, the bunch length $\sigma_{z}$ (Fig.8) increases rapidly during acceleration due to the velocity–energy correlation, as faster electrons at the head outpace slower ones at the tail; beyond the cavity, the expansion continues more gradually under residual space‐charge forces.
Figure 7: Simulated mean kinetic energy of the electron bunch along the beamline. The beam rapidly reaches relativistic energies in the RF photogun, exiting with mean energy $\overline{W_{\mathrm{kin}}}=1.9\ \mathrm{MeV}$.
The simulated energy evolution is defined by the presently available RF power. Whereas the gun can in principle deliver up to 5MeV5\ $\mathrm${MeV} electrons at its design RF-power limit of 6MW6\ $\mathrm${MW}, only 3MW3\ $\mathrm${MW} are available in the current hardware configuration. Consequently, the accelerating gradient is reduced to $\approx 45\ \mathrm{MV/m}$, resulting in a final kinetic energy of $\approx 1.9\ \mathrm{MeV}$ in Fig. 7. The reduced gradient also increases the relative importance of space-charge forces during the first few millimeters after the emission.
Figure 8: Simulated RMS bunch length $\sigma_{z}$ along the beamline. Rapid longitudinal expansion occurs during RF photogun acceleration, followed by gradual growth from energy spread and residual space-charge effects.
Transversely, the beam expands symmetrically in $x$ and $y$ (Fig. 9(a)) from 2 mm to nearly 25 mm in the absence of external focusing. At the present low bunch charge (0.63pC0.63\ $\mathrm${pC}) and relatively large emission area, the space‐charge forces are moderate and contribute mainly to the divergence of the transverse envelope. The evolution of the normalized emittance, however, is dominated by correlated distortions from the transverse RF fields of the 1.5-cell $\pi$-mode gun, rather than by space charge. As shown in Fig. 9(b), the emittance remains near $1.5\ \pi\cdot\mathrm{mm}\cdot\mathrm{mrad}$ at the cavity exit but exhibits early oscillations characteristic of RF-induced correlations. These oscillations become particularly relevant when a solenoidal field is introduced, since the RF–solenoid coupling modifies their cancellation. A slight xxyy asymmetry likely originates from numerical sampling noise. Without external focusing, the beam size diverges rapidly and RF-induced emittance growth cannot be mitigated, motivating the use of a compensation solenoid in the following section.
Figure 9: Simulated beam dynamics without emittance compensation. (a): Simulated RMS transverse beam sizes $\sigma_{x}$ and $\sigma_{y}$ along the beamline, showing symmetric growth from 2mm2\ $\mathrm${mm} to nearly 25mm25\ $\mathrm${mm}. (b): Simulated evolution of the normalized transverse emittance along the beamline. After initial RF-induced oscillations inside the 1.5-cell gun, the emittance remains approximately constant at the cavity exit and subsequently increases due to nonlinear RF-field–induced transverse phase-space distortions.
Because the space-charge effects are weak in this regime, they can be neglected in the subsequent analysis. This is particularly important for the quantum-mechanical considerations in Sec. IV, since the small Coulomb forces do not significantly perturb the phase structure or the orbital-angular-momentum content of the evolving electron wave packet.
III.2 Emittance Compensation Using a Solenoidal Lens
A solenoidal lens is installed near the cathode (Fig. 4) to counteract early nonlinear space‐charge effects. The on‐axis field profile used in the simulations (Fig. 5) was optimized via iterative scans, yielding a peak of $B_{z}=125$ mT at $z=85$ mm downstream of the photocathode.
Fig. 10(a) shows the resulting RMS transverse beam size. Immediately after the RF cavities, the solenoid collimates the envelope, stabilizing $\sigma_{x,y}$ along the beamline. The normalized emittance (Fig. 10(b)) initially rises due to the non-zero solenoidal field at the cathode, which imparts extrinsic angular momentum onto the beam. In addition, the combination of the solenoidal field with the transverse RF fields of the 1.5-cell $\pi$-mode gun modifies the evolution of the transverse phase space and suppresses the natural cancellation of RF-induced correlations that occurs in the zero-solenoid case [Floettmann15].
Figure 10: Simulated beam dynamics using the emittance compensation solenoid. (a): Simulated RMS transverse beam size along the beamline with the compensation solenoid. The size stabilizes soon after the RF cavities, indicating effective collimation. (b): Simulated normalized transverse emittance along the beamline with solenoid compensation. Stabilizes near $3.2\ \pi\cdot\mathrm{mm}\cdot\mathrm{mrad}$ downstream.
Parametric scans further validate the robustness of the chosen settings: Fig. 11(a) confirms the optimum solenoid peak at z=85z=85\, mm (fixed $B_{z}=125\,$ mT), and Fig. 11(b) shows that $B_{z}=125\,$ mT yields the minimum final emittance when positioned at this location. These settings provide a stable and experimentally accessible working point for the injector under the present operating conditions.
Figure 11: Final normalized transverse emittance versus solenoid parameters: (a) dependence on the solenoid peak position for a fixed field of 125125\,mT; (b) dependence on the peak solenoid field strength with the solenoid fixed at z=85z=85\,$m$. In both cases the minimum emittance is achieved for a peak field of 125125\,mT at z=85z=85\,$m$.
IV Quantum Beam Dynamics
The classical beam dynamics simulations presented above demonstrates that the photoinjector can operate in a regime of low bunch charge and correspondingly weak space-charge forces, yielding a stable, low-emittance beam with minimal RF-induced phase-space distortions. This establishes that the test bench is capable of producing high-quality electron wave packets under realistic operating conditions. In this regime, it becomes essential to evaluate whether structured quantum states – specifically, vortex electrons – can be generated and accelerated while preserving their spatial coherence and OAM content. Motivated by the experimentally accessible values of orbital angular momentum that can be imparted by our UV vortex optics, we now turn to the quantum description of single-electron Laguerre–Gaussian states in the RF fields of the photoinjector.
IV.1 Vortex Electrons
In general, a vortex electron beam is a free-electron state whose wave function can be represented in terms of an LG mode, an exact non-stationary solution of the Schrödinger equation and a close analogue of optical vortex modes [Allen92]. Its transverse wave function can be written as [Karlovets21]
$\begin{split}\psi_{\ell,n}(\bm{\rho},t)&=N\frac{\rho^{|\ell|}}{(\sigma_{\perp}(t))^{|\ell|+1}}L_{n}^{|\ell|}\left(\frac{\rho^{2}}{(\sigma_{\perp}(t))^{2}}\right)\\ &\quad\times\exp[i\ell\phi_{r}-i(2n+|\ell|+1)\arctan\left(\frac{t}{t_{d}}\right)\\ \hfil&\qquad-\frac{\rho^{2}}{2(\sigma_{\perp}(t))^{2}}\left(1-i\frac{t}{t_{d}}\right)\Biggr]\end{split}$ (1)
where the normalization constant $N$ is chosen such that
$\int d^{2}x\,|\psi_{\ell,n}(\bm{\rho},t)|^{2}=1.$ (2)
Here, $\ell=0,\pm 1,\pm 2,$..., is the OAM in units of $\hbar$, n=0,1,2,,n=0,1,2,..., is the radial quantum number, $L_{n}^{\ell}$ denotes the generalized Laguerre polynomial, $\sigma_{\perp}(t)$ is the transverse beam size of the fundamental LG mode: for $n=\ell=0$, and $t_{d}$ is the characteristic diffraction time. It is important to note that the azimuthal phase factor $\exp(i\ell\phi_{r})$ makes the state an eigenfunction of the OAM operator,
$\hat{L}_{z}\,\psi_{\ell,n}=-i\hbar\frac{\partial}{\partial\phi_{r}}\psi_{\ell,n}=\ell\hbar\,\psi_{\ell,n},$
and it is the defining feature of vortex beams: it produces the characteristic helical phase front and ensures that each electron carries an orbital momentum projection $\ell\hbar$ with respect to the propagation axis.
One of the mechanisms for generating the vortex electrons that can be realized at an accelerator facility in an RF photoinjector is photoemission driven by structured light. Theory shows that an atom placed on the axis of an optical vortex transfers a well-defined OAM projection to the emitted electron, and this behaviour persists for ensembles [Pavlov24, Kazinski25]. Alternative schemes based on electron diffraction [Uchida10, Verbeeck10, Blackburn14] are impractical in such setups.
IV.2 Quantum Evolution of a Single-Electron Wave Packet in the RF Field
To analyze the generation of relativistic vortex electrons carrying OAM, we follow the quantum evolution of a single-electron wave packet during acceleration. As a convenient observable we consider the mean-squared transverse radius $\langle\rho^{2}\rangle(\langle z\rangle)$, which for an LG wave packet is related to the transverse rms width via [Karlovets21]
$\langle\rho^{2}\rangle(\langle z\rangle)=\sigma_{\perp}^{2}(\langle z\rangle)\,(2n+|\ell|+1).$ (3)
Throughout this section, the angular brackets denote quantum expectation values over the single-electron state. We compare three representative cases::
  1. 1.
    Free-space propagation over the RF photogun length;
  2. 2.
    Acceleration in a uniform DC electric field, corresponding to a simplified DC-gun model;
  3. 3.
    Acceleration in the realistic standing-wave RF field of the photoinjector.
The general law of transverse quantum spreading for free (field-less) wave packets [Karlovets21] can be written either as a function of the mean propagation distance $\langle z\rangle$ or of the corresponding time of flight $t(\langle z\rangle)$:
$\langle\rho^{2}\rangle(\langle z\rangle)=\langle\rho^{2}\rangle(0)\left(1+\frac{\langle z\rangle^{2}}{z_{R}^{2}}\right)=\langle\rho^{2}\rangle(0)\left(1+\frac{t^{2}(\langle z\rangle)}{t_{d}^{2}}\right).$ (4)
Here $z_{R}$ is the Rayleigh length and $t_{d}$ is the diffraction time [Karlovets21],
$z_{R}=\frac{2\pi}{M}\frac{\langle\rho^{2}\rangle(0)}{\lambda_{\mathrm{dB}}},\qquad t_{d}=\frac{m\,\langle\rho^{2}\rangle(0)}{\hbar M},$ (5)
where $M=2n+|\ell|+1$ is the beam-quality factor of the packet and $\lambda_{\mathrm{dB}}$ is the de Broglie wavelength. Note that Eqs. (4) and (5) depend on the quantum numbers $(n,\ell)$ only through the combination $M$ and the initial second moment $\langle\rho^{2}\rangle(0)$; therefore the same transverse spreading law applies to any coherent single-electron state, not only to the LG modes, provided that they share the same initial rms transverse radius.
In the far-field regime, $\langle z\rangle\gg z_{R}$, Eq. (4) reduces to a simple relation, often referred to as a van Cittert–Zernike–type theorem [Karlovets21], between the measured packet width at the detector and the initial width near the source,
$\sqrt{\langle\rho^{2}\rangle(0)}=\frac{\langle z\rangle\,\lambda_{\mathrm{dB}}}{2\pi\,\sqrt{\langle\rho^{2}\rangle(\langle z\rangle)}}\,M.$ (6)
In the intermediate (Fresnel) regime the transverse width obeys
$\begin{split}\sqrt{\langle\rho^{2}\rangle(\langle z\rangle)}={}&\frac{\langle z\rangle\,\lambda_{\mathrm{dB}}}{2\pi\,\sqrt{\langle\rho^{2}\rangle(0)}}\,M\\ &\times\left[1+\frac{1}{2}\left(\frac{2\pi\,\langle\rho^{2}\rangle(0)}{\langle z\rangle\,\lambda_{\mathrm{dB}}}\,\frac{1}{M}\right)^{2}\right].\end{split}$ (7)
For an initial rms transverse radius $\sqrt{\langle\rho^{2}\rangle(0)}=1\ \mathrm{nm}$, typical for an electron packet near the emission region [Karlovets21, Ehberger15], with $n=3$ and $\ell=1$ (OAM $\ell\hbar$) and a drift length of 30cm30\ $\mathrm${cm}, Eq. (4) predicts free-space spreading to an rms transverse width of about 0.8m0.8\ $\mathrm${m} for an initial kinetic energy of 0.3eV0.3\ $\mathrm${eV}. Next, we include acceleration.
Uniform DC acceleration.
In a static electric field $\mathbf{E}=E\,\hat{\mathbf{z}}$ with a scalar potential $A^{0}=-Ez$ and $\mathbf{A}=0$, the Hamiltonian separates as
$\hat{H}=\frac{p_{x}^{2}+p_{y}^{2}}{2m}+\left(\frac{p_{z}^{2}}{2m}+qEz\right)\equiv\hat{H}_{\perp}+\hat{H}_{z},$ (8)
so that $\hat{H}_{\perp},\hat{H}_{z}]=[\rho^{2},\hat{H}_{z}]=0$. The transverse quantum evolution is therefore identical to free space; the only modification arises through the shortened time of flight to a given longitudinal coordinate $z$. Using $\varepsilon(z)=\varepsilon_{0}+qEz$ and $p(z)=\sqrt{\varepsilon(z)^{2}-m^{2}c^{4}}/c$, one finds
$t(z)=\frac{p(z)-p_{0}}{qE},$ (9)
where for acceleration of electrons along +z+z one has $E<0$ and hence $qE>0$, so that $t(z)>0$. The transverse second moment at a fixed plane $z$ is then obtained by evaluating Eq. (4) at this $t(z)$. For typical DC-gun this results in a $\sim 550$-fold reduction of the final transverse width compared to free-space propagation (Fig. 12), in good agreement with recent relativistic calculations [z6j7-grs2].
Figure 12: Transverse intensity distributions of the LG wave packets ($n=3$, $\ell=2,4,10,64$) with the initial rms transverse radius $\sqrt{\langle\rho^{2}\rangle(0)}=1\ \mathrm{nm}$ and initial kinetic energy $K_{0}=0.39\ \mathrm{eV}$ after $30~\mathrm{cm}$ of propagation. Left column: free-space propagation. Right column: acceleration in a uniform DC field with $E=-2~\mathrm{MV/m}$. The normalized maps $|\psi(x,y)|^{2}$ at the observation plane illustrate the strong suppression of transverse spreading under DC acceleration while preserving the characteristic ring structure of the LG modes.
RF standing-wave acceleration.
In the S-band photoinjector electrons are accelerated in a standing-wave RF field. Along the axis, where the packet propagates, the longitudinal electric field can be written as
$E_{z}(z)=E_{z0}\,\sin\!\left(\omega\frac{z}{v}+\phi\right),$ (10)
where $E_{z0}$ is the peak amplitude, $\omega$ is the RF angular frequency, and $\phi$ is the injection phase. The corresponding kinetic-energy gain over an accelerating gap of length $d$ is
$\Delta E_{\rm kin}=eE_{z0}\int_{-d/2}^{d/2}\!\cos\!\left(\omega\frac{z}{v}+\phi\right)\,dz,$ (11)
which accounts for the phase-dependent acceleration experienced by particles with finite longitudinal extent and underlies the development of a velocity–energy correlation during bunching. The on-axis energy profile $K(z)$ used below is obtained from the simulated field map.
Hamiltonian structure. In the TM010 standing-wave mode the on-axis electromagnetic field is predominantly longitudinal. Near the axis the field admits the expansions
$E_{z}(\rho,z,t)$ $=E_{z}(0,z,t)+\mathcal{O}\!\left(\rho^{2}\right),$ (12)
$B_{\varphi}(\rho,z,t)$ $=\mathcal{O}\!\left(\rho/R_{\mathrm{cav}}\right),$ (13)
where $R_{\mathrm{cav}}$ is the cavity radius. For a nanometer-scale packet one has $\rho/R_{\mathrm{cav}}\lesssim 10^{-8}$, so that $|qvB_{\varphi}|/|qE_{z}|\ll 10^{-7}$ and transverse magnetic forces are negligible.
In a gauge with $A^{0}=0$ and a purely longitudinal vector potential $A_{z}(z,t)$ satisfying $E_{z}=-\partial_{t}A_{z}$, the single-particle Hamiltonian takes the form
$\hat{H}(t)=\frac{p_{x}^{2}+p_{y}^{2}}{2m}+\frac{\bigl[p_{z}-qA_{z}(z,t)\bigr]^{2}}{2m}\equiv\hat{H}_{\perp}+\hat{H}_{z}(t),$ (14)
and, because $A_{z}$ depends only on $z$ and $t$,
$\hat{H}_{\perp},\hat{H}_{z}(t)]=0,\qquad[\rho^{2},\hat{H}_{z}(t)]=0,$ (15)
up to corrections of order $(\rho/R_{\mathrm{cav}})^{2}$. Thus the transverse dynamics remains effectively free, as in the DC case, and the RF field enters only through the longitudinal kinematics.
Longitudinal evolution and time-of-flight. The kinetic energy satisfies
$K(z)=K_{0}+\Delta E_{\rm kin}(z),\qquad\gamma(z)=1+\frac{K(z)}{mc^{2}},$ (16)
and the local velocity is $\beta(z)=\sqrt{1-\gamma(z)^{-2}}$. The flight time to a plane at $z$ follows from
$t(z)=\int_{0}^{z}\frac{dz^{\prime}}{\beta(z^{\prime})\,c}.$ (17)
For the photoinjector parameters used in this work, the resulting relativistic velocity profile reduces the time of flight to the exit plane by almost three orders of magnitude compared to free-space propagation. In the regime $t\gg t_{d}$, where the transverse width obeys $\sigma_{\perp}(t)\propto t$, this kinematic shortening leads to an $\sim 900$-fold suppression of the transverse rms size compared to the free electron at $z=30~\mathrm{cm}$, as shown in Fig. 13. Importantly, similar suppression of quantum spreading persists for all other parameter sets explored in this work (not shown), indicating that strong reduction of transverse spreading is a robust feature of relativistic acceleration rather than a special property of the specific cases illustrated in Figs. 1312.
Figure 13: Transverse intensity distributions of the LG wave packets ($n=3$, $\ell=2,4,10,64$) with the initial rms transverse radius $\sqrt{\langle\rho^{2}\rangle(0)}=1\ \mathrm{nm}$ and initial kinetic energy $K_{0}=0.39\ \mathrm{eV}$ at $z=30~\mathrm{cm}$ for free-space propagation (left) and acceleration in the standing-wave RF field of the S-band photogun (right). The normalized maps $|\psi(x,y)|^{2}$ demonstrate the strong suppression of transverse spreading during RF acceleration while preserving the characteristic LG ring structure for all OAM values.
These propagation laws, derived for single-particle packets without accounting for collective effects, can be directly applied to structured vortex beams. This provides guidance for preserving spatial coherence and OAM content during beam transport and acceleration. By understanding and controlling this fundamental quantum spreading, photoinjector designs can be optimized to maintain the quality and characteristic properties of the vortex beams throughout the early acceleration stages.
V Conclusions
We have performed classical and quantum beam-dynamics simulations of the S-band RF photoinjector test bench now being commissioned at JINR. Using realistic electromagnetic field maps and ASTRA tracking, we have showed that the injector can operate in a stable ultralow-charge regime ($Q\approx 0.63$ pC), where space-charge forces are weak and the emittance evolution is dominated by RF-induced correlations. Under these conditions, a Carlsten-type emittance-compensation scheme was implemented and optimized. Although the solenoid does not reduce the emittance below its intrinsic value, it suppresses RF-driven phase-space distortions and provides a reproducible operating point with a final normalized emittance of $\approx 3.2\ \pi\cdot\mathrm{mm}\cdot\mathrm{mrad}$.
The low-charge operating regime has enabled a quantum-mechanical analysis of single-electron states. We have studied the evolution of Laguerre–Gaussian wave packets with the realistic OAM values under three propagation scenarios: free drift, acceleration in a uniform DC field, and acceleration in the standing-wave RF field of the S-band gun. In both accelerating cases, the rapid growth of electron momentum has substantially suppressed transverse spreading compared to free-space propagation, while preserving the characteristic radial structure of the OAM-carrying states.
Overall, the simulations show that the present injector configuration satisfies the beam-quality requirements for experiments on the photoemission-driven generation and acceleration of vortex electrons in the MeV energy range. The photoinjector has been assembled, and the first photoemitted electron beam has been produced. Commissioning of the RF gun is ongoing, with work underway to reach the design input power of 6 MW. In parallel, a phase-locking system for synchronizing the UV laser pulses with the RF field is being developed. All preparations for experiments on photoemission-based generation of relativistic vortex electrons have been completed: diffractive optical elements for producing ultraviolet twisted light beams with orbital angular momentum up to $\ell=64\hbar$ have been fabricated and tested. Future studies will extend the present modelling by including collective interactions and decoherence mechanisms to obtain a complete description of structured electron beams under realistic experimental conditions. The check of the electron beams vorticity can be made by diffraction on a triangle apperture, as has been argued in Ref. [Maksimov25].
Acknowledgments
The authors express their gratitude to A.V. Vukolov, M.V. Shevelev, and G.K. Sizykh for their significant assistance in developing the beam emittance diagnostics. We also thank E.A. Yakushev, V.V. Glagolev, D. V. Naumov, A.S. Zhemchugov, and the management of the Laboratory of Nuclear Problems for supporting the project on behalf of JINR. Special thanks are due to K.V. Cherepanov and N.E. Sheremet for their help with quantum dynamics calculations, and to S.S. Baturin for valuable consultations.
This study was supported by the Russian Science Foundation (Project No. 23-62-10026) [RSF].